OpenStudy (anonymous):

Solve and show work. 5 over 7 minus one half

OpenStudy (lollylau):

5/7-1/2 =5*2-1*7/7*2 =3/14

OpenStudy (lollylau):

lol 'pi'

OpenStudy (anonymous):

5/7 - 1/2 = 3/14

OpenStudy (lollylau):

it says show work =.="

OpenStudy (anonymous):

No, pi is 3.14, this is 3/14.

OpenStudy (lollylau):

"pi" quotes

OpenStudy (anonymous):

thanks i know how you got that answer now